# Neuroplastic Expansion in Deep Reinforcement Learning

### v1

@@ Results
In this section, we conduct an additional ablation study to investigate the effectiveness of neuron consolidation in Neuroplastic Expansion, while the importance of elastic neural generation and dormant neural pruning has been demonstrated in Figure 3. We also examine the experience review (ER) module and the impact of our method on the policy and value networks.
@@
Figure 7: Ablation study: (a) shows the validity verification of ER, while (b) to (d) detail the impact of NE on the actor and critic from the perspectives of plasticity and performance.
As illustrated in Figure 7a, the experience review technique enhances performance stability across different seeds. In contrast, the absence of this critical component leaves the model vulnerable to potential policy collapse in subsequent stages (albeit slighter compared to Reset), demonstrating the critical role of reviewing historical knowledge in maintaining policy stability and preventing the loss of early valuable information as discussed in Section 4.2.
Figure 7b illustrates the comparison result of our method and its variants, isolating the effects of NE on the actor and critic networks separately.
We observe that NE is particularly crucial for the critic network. Removing this from the critic leads to significant performance degradation and significant loss of plasticity compared to the full version of NE (Figure 7c,7d); whereas limiting the expansion to only the critic results in a less pronounced drop in performance. This asymmetry in impact may be attributed to the more challenging task faced by the critic, as it needs to provide accurate value estimates for different state-action pairs (Meng et al., 2021) considering a non-stationary data distribution. In addition, the critic also provides guidance signals to actor optimization, which requires high level of adaptability (Fujimoto et al., 2024).
5.2 Continual adaptation
In this section, we investigate the continual ability to learn and adapt to changing environments, which is a key ability of deep RL agents (Abbas et al., 2023, Elsayed and Mahmood, 2024).
@@ 5.2 Continual adaptation
[…]
Figure 8: Plasticity measures, i.e., activated neuron ratio, id recorded for three agents with different settings during long term cyclic training.
@@ Results
The performance of each method for a single environment when repeatedly learning on a sequence of environments as described above is summarized in Figure 9. We aim to evaluate (i) the efficiency of the policy in swiftly learning the subsequent task after mastering the previous one within a single cycle, (ii) the agent’s ability to retain the benefits of initial learning for the same task across multiple cycles. As shown, Resetting, which involves initializing the parameters after each task, is currently deemed the most effective approach.
According to the row in Figure 9, NE adeptly handles each task without being influenced by previous training. It significantly outperforms the original TD3, emphasizing that standard RL algorithms do suffer from catastrophic forgetting in the continual learning setting. We attribute this success to NE’s capacity to introduce real-time topology adjustments for storing new knowledge, thereby ensuring continuous learning capabilities. Remarkably, in certain environments, such as Ant, NE-TD3 can retain previous knowledge and seamlessly continue learning. This could be attributed to the experience review module and the preservation of old neurons in the topology, which store pertinent information crucial for successful continual learning. When considering the column-wise comparisons, our method performs comparably to reset in the majority of tasks, yet significantly outshines reset in the challenging Humanoid environment. The comparison in terms of active neurons for each baseline is shown in Figure 8, which illustrates that even in the long training schedule, NE can effectively alleviate neuron dormancy while efficiently learning, and it even performs better than resetting all parameters during the early training period.
(a) First task: HalfCheetah
(b) Second task: Humanoid
(c) Third task:Ant
(d) Fourth task:Hopper
5.3 Versatility
In this section, we demonstrate the generality of our proposed method by applying it to another popular deep RL method and evaluating its performance on more complex tasks with image input.
Experimental Setup
We evaluate four image-based motion control tasks from the DeepMind Control Suite (DMC) (Tassa et al., 2018b): Reacher hard, Reacher easy, Walker walk and Cartpole Swingup Sparse, to demostrate the generalization of NE for deep RL algorithms and its robustness across diverse tasks. We consider DrQ (Yarats et al., 2022), a well-known variant of SAC (Haarnoja et al., 2018) specifically designed for processing image inputs, as the backbone algorithm. We use the same baselines from Section 5.1 for comparison.
Figure 10: First row shows the performance of five methods. Second row corresponds to the percentage of activated neurons. All experiments run with seven independent seeds
Results
First row of Figure 10 shows that NE can be effectively combined with DrQ to achieve stable performance on image input tasks, outperforming other baselines in all four scenarios and demonstrating its versatility across different algorithms. In addition, the second row of Figure 10 (activated neuron ratio) indicates that, although EN cannot reach the peak value set by Reset, it can better delay the dormancy of neurons compared to other warm mitigation startegies and ensure stable performance.
@@ 6 Conclusion
Inspired by the cortical growth in biological agents that triggers neuronal topology expansion to enhance plasticity and adaptability to new situations, we present a novel perspective for deep RL agents to mitigate the loss of plasticity with a dynamic growing network. Building on this insight, we introduce a comprehensive topology growth framework, Neuroplastic Expansion (NE), which addresses both growth and pruning perspectives to maximize the performance benefits of topology expansion. NE adds high-quality elastic neurons and connections based on the gradient to improve the network’s learning ability; prunes dormant neurons in a timely manner for better utilizing network capacity, and returns them to the candidate set to maintain the agent’s plasticity, making dormant neurons reusable. NE outperforms previous methods which aimed at alleviating plasticity loss on diverse tasks and demonstrates stable, outstanding performance on plasticity metrics. Interesting future directions include the exploration of adaptive architectures that dynamically adjust their capacity based on the agent’s performance or task complexity, and further minimizing computational overhead. Overall, despite the acknowledged limitations, our research provides practical insights into enhancing the plasticity of RL agents. We hope our findings pave the way for diverse future explorations, potentially leading to more sample- and parameter-efficient, as well as adaptable, RL algorithms.
Appendix A Related Preliminaries
@@
$L_{Q}(\theta_{i})=\mathbb{E}_{s,a,s^{\prime}}\big{[}\big{(}y-Q_{\theta_{i}}(s,a)\big{)}^{2}\big{]}\text{ for }\forall i\in\{1,2\}.$ (3)
Where $y=r+\gamma\min\limits_{i=1,2}{Q_{\bar{\theta}_{i}}}(s^{\prime},\pi_{\bar{\phi}}(s^{\prime}))$, $\bar{\phi}$ denotes the target network parameters. The actor $\pi_{\phi}$ is updated according to the Deterministic Policy Gradient (Fujimoto et al., 2018):
$\nabla_{\phi}J(\phi)=\mathbb{E}_{s}\big{[}\nabla_{\pi_{\phi}(s)}Q_{\theta_{1}}(s,\pi_{\phi}(s))\nabla_{\phi}\pi_{\phi}(s)\big{]}.$ (4)
At each training step $t$, the agent first randomly samples a batch of state transitions from the buffer and calculates the loss gradient as described above. It is worth noting that the proposed method in our paper is a plug-in and does not affect the training process of the reinforcement learning algorithm itself. For this reason, the training details of the reinforcement learning algorithm are abbreviated in the main text.
@@
DrQ is an algorithm based on maximum entropy reinforcement learning. Its backbone network and training method are the same as SAC (Haarnoja et al., 2018). The soft value function is trained to minimize the squared residual error:
$J_{V}(\phi)=\mathbb{E}_{(s,a)\sim D}{\frac{1}{2}(V_{\phi}(s)-Q_{\theta}(s,a)+\log\pi_{\psi}(a|s))}$ (5)
where $a$ are sampled according to the current policy, instead of the replay buffer. The soft Q-function parameters can be trained to minimize the soft Bellman residual: $J_{Q}(\theta)=\mathbb{E}_{(s,a)\sim D}[\frac{1}{2}(Q_{\theta}(s,a)-\hat{Q}(s,a))^{2}]$. Then the policy network can be learned by minimizing the expected KL divergence
$J_{\pi}(\psi)=\mathbb{E}_{s\sim D}\big{[}D_{KL}\big{(}\pi_{\phi}(\cdot|s)||\frac{\exp{Q(s)}}{Z(s)}\big{)}\big{]}$ (6)
Appendix B Experimental Details
B.1 Network Structure
@@
The normal size of the TD3 network we used is the official architecture, and the detailed setting is shown in Tab 1. Besides, the network we used for Scale-up experiments incrementally adds the depth of both the actor and critic. For The 1.8x agent, we add one MLP as a hidden layer in two critic networks in front of the output head and one MLP layer as the embedding layer at the front of the actor. And the hidden dim is also increased to 256 (double). And so on, we follow the above approach to grow the model by an equal amount each time.
@@
Fully Connected (128, action dim) (128, 1)
@@ B.2 Hyperparameter
Our codes are implemented with Python 3.8 and Torch 1.12.1. All experiments were run on NVIDIA GeForce GTX 3090 GPUs. Each single training trial ranges from 6 hours to 21 hours, depending on the algorithms and environments, e.g. DrQ spends more time than TD3 to handle the image input and DMC needs more time than OpenAI mujoco for rendering. Our TD3 is implemented with reference to github.com/sfujim/TD3 (TD3 source-code). The hyper-parameters for TD3 are presented in Table 2. Notably, for all OpenAI mujoco experiments, we use the raw state and reward from the environment and no normalization or scaling is used. An exploration noise sampled from $N(0,0.1)$ (Lillicrap, 2015) is added to all baseline methods when selecting an action. The discounted factor is $0.99$ and we use Adam Optimizer (Kingma, 2014) for all algorithms. Tab.2 shows the common hyperparameters of TD3 used in all our experiments. And for NE, following the related work (Elsayed et al., 2024), we use weight Clip on dormant neurons which are already reset and pruned from the current topology. Notably, We add layer normalization after the hidden layer of both critic and actor to construct the comparison baseline LN-TD3 To construct the comparison baseline LN-TD3, but in all other methods, no normalization mechanism is introduced to ensure the fairness of comparison. We reproduce plasticity Injection according to their official paper (Nikishin et al., 2024). And use the official code of ReDo in our comparison under the default setting https://github.com/google/dopamine/tree/master/dopamine/labs/redo. For Reset-TD3, after a lot of empirical debugging, we found that parameter reset every $90,000$ step is a generally better setting, and fixed it as a hyperparameter. We use seeds 0 to 7 for all the experiments. And if you want to reproduce the HalfCheeetah learning curve, we suggest you use seeds 0 5. For reset-TD3, reset is used every $90,000$ steps
@@
Representation Model Learning Rate None None None None None None
@@
Table 2: A comparison of common hyperparameter choices of algorithms. We use None to denote the ’not applicable’ situation.

### v3

@@ 4.3 Ablation Study
In this section, we conduct an additional ablation study to investigate the effectiveness of neuron consolidation in Neuroplastic Expansion, while the importance of elastic neural generation and dormant neuron pruning has already been demonstrated in Figure 3. Additionally, we examine the experience review (ER) module and assess the impact of our method on the policy and value networks.
@@
Figure 9: Ablation study: (a) presents the validity verification of ER, while (b) to (d) illustrate the impact of NE on the actor and critic from the perspectives of plasticity and performance.
As illustrated in Figure 9a, the results on HalfCheetah show ER improves the training stability, particularly in the later stage. Additionally, we select 4 tasks with image input and sparse reward setting to further verify the effectiveness of ER. The results in Appendix F.4 indicate that when assisted with ER, our training framework consistently stabilizes the convergent policy at a high level across different seeds (i.e., lower standard deviation). In contrast, the absence of this critical component makes the model more vulnerable to policy collapse in later stages (albeit to a lesser extend than Reset). This demonstrates the importance of reviewing historical knowledge in maintaining policy stability and preventing the loss of early valuable information, as discussed in Section 3.2.
Figure 9b illustrates a comparative analysis of our method and its variants, isolating the effects of NE on the actor and critic networks separately. We observe that NE plays a particularly crucial role in the critic network. Removing NE from the critic results in significant performance degradation and substantial plasticity loss compared to the full version of NE (Figure 9c, 9d). In contrast, limiting expansion to only the critic leads to a less pronounced drop in performance. This asymmetry in impact may be attributed to the greater challenge faced by the critic, as it must provide accurate value estimates for various state-action pairs (Meng et al., 2021) under a non-stationary data distribution. In addition, the critic provides guidance signals for actor optimization, which requires a high level of adaptability (Fujimoto et al., 2024).
4.4 Versatility
In this section, we demonstrate the generality of our proposed method by applying it to another popular deep RL algorithm and evaluating its performance on more complex tasks with image inputs.
@@ 4.4 Versatility
We evaluate four image-based motion control tasks from the DeepMind Control Suite (DMC) (Tassa et al., 2018b): Reacher Hard, Reacher Easy, Walker Walk, and Cartpole Swingup Sparse, to demonstrate the generalization of NE across deep RL algorithms and its robustness across diverse tasks. For the backbone algorithm, we use DrQ (Yarats et al., 2022), a well-known variant of SAC (Haarnoja et al., 2018) specifically designed for processing image inputs. We compare our approach against the same baselines used in Section 4.1. The weight clipping parameter $\kappa$ is set to 2.
Figure 10: The first row shows the performance of five methods. The second row corresponds to the percentage of activated neurons. All experiments were run with seven independent seeds.
@@ Results
The first row of Figure 10 shows that NE can be effectively combined with DrQ to achieve stable performance on image-input tasks, outperforming all baselines and demonstrating its versatility across different algorithms. Additionally, the second row of Figure 10 (activated neuron ratio) indicates that although NE does not reach the peak activation levels of Reset, it delays neuron dormancy more effectively than other warm mitigation strategies, ensuring stable performance.
@@ 5 Conclusion
Inspired by the cortical growth in biological agents that triggers neuronal topology expansion to enhance plasticity and adaptability to new situations, we introduce a novel perspective for deep RL agents to mitigate the loss of plasticity through a dynamic growing network. Building on this insight, we present a comprehensive topology growth framework, Neuroplastic Expansion (NE), which addresses both growth and pruning mechanisms to maximize the performance benefits of topology expansion. NE adds high-quality elastic neurons and connections based on the gradient signals to improve the network’s learning ability; prunes dormant neurons in a timely manner for better utilizing network capacity, and returns them to the candidate set to maintain the agent’s plasticity, making dormant neurons reusable. NE outperforms previous methods that aimed at alleviating plasticity loss on diverse tasks and demonstrates stable, outstanding performance on plasticity metrics. Future directions include exploring adaptive architectures that dynamically adjust their capacity based on the agent’s performance or task complexity, as well as further reducing computational overhead. Overall, despite its acknowledged limitations, NE provides practical insights into enhancing the plasticity of deep RL agents. We hope our findings pave the way for future research, potentially leading to more sample-efficient and adaptable deep RL algorithms.
Acknowledgment
This work is supported by the National Natural Science Foundation of China 62406266. We would also like to thank the Python community (Van Rossum & Drake Jr, 1995; Oliphant, 2007) for developing tools that enabled this work, including NumPy (Harris et al., 2020) and Matplotlib (Hunter, 2007).
Appendix A Related Work
Plasticity loss in RL
Recent studies have found that agents guided by the non-stationary objective characteristic of RL suffer catastrophic degradation (Lyle et al., 2023) and ultimately overfit early experiences, which limits their continuous learning ability. This phenomenon is known as plasticity loss. Behaviorally, the loss of plasticity during training can manifest as, for instance, an inability to obtain an effective gradient to guide policy adaptation in fine-tuning (Dohare et al., 2024), limiting potential transferability and making it difficult to adapt to new sampling and unseen data (refer to as primacy bias (Nikishin et al., 2022)). While the research on this topic is still in its early stages, several techniques have been demonstrated to mitigate the loss of plasticity in deep RL. Resetting global networks or dormant neurons  (Nikishin et al., 2022; Sokar et al., 2023) and modifying activation functions (Abbas et al., 2023) are observed to mitigate plasticity loss. Lyle et al. (2024b) empirically finds that network normalization, particularly layer regularization (LN), can maintain plasticity. Nikishin et al. (2024) introduced copies of random heads for injecting plasticity to improve continual learning ability. PLASTIC (Lee et al., 2023) divides plasticity into input plasticity and label plasticity, addressing them separately. NaP (Lyle et al., 2024a) alleviates the over-fitting issue from the perspective of the association between regularization and learning rate. In pixel-based deep RL, data augmentation (Ma et al., 2023) and batch size reduction (Obando Ceron et al., 2023) have been analyzed for their effectiveness in reducing plasticity loss. Recent studies have found that mitigating gradient starvation (Dohare et al., 2024) and constraining deviations from initial weights (Kumar et al., 2023; Lewandowski et al., 2024) can also be effective in mitigating plasticity loss. In supervised learning, Lee et al. (2024a) proposed to use a combination of fast and slow update networks to balance between fast, fleeting adaptation and slow, steady generalization in supervised learning. Ash & Adams (2020); Berariu et al. (2021) found that both warm initialize and pre-train can better adapt to increasing amounts of data.
Pruning in RL
Sokar et al. (2021) showed that training the deep RL policy with a changing topology is difficult due to training instability. Since then, this challenging topic has been well-studied. Policy Pruning and Shrinking (PoPs) (Livne & Cohen, 2020) obtain a sparse deep RL agent with iterative parameter pruning. Graesser et al. (2022); Tan et al. (2022); Lee et al. (2024b) attempt to train a sparse neural network from scratch without pre-training a dense teacher. Existing methods mainly focus on obtaining a smaller topology to improve training efficiency. However, our paper explores whether growing the topology from small to large can mitigate the plasticity loss in deep RL agents rather than training with the maximum number of parameters. In deep RL, the model size commonly utilized is typically small and may encounter policy collapse after scaling up (Schwarzer et al., 2023; Ceron et al., 2024a). Thus, sparse training techniques that have proven beneficial for significantly reducing training costs in fully supervised domains may not yield similar advantages in deep RL in terms of computation cost reduction. Instead, we provide a new perspective and focus on making deep RL agents better use of the growing network size to further improve their learning ability.
Appendix B Related Preliminaries
@@
$L_{Q}(\theta_{i})=\mathbb{E}_{s,a,s^{\prime}}\big{[}(y-Q_{\theta_{i}}(s,a))^{2}\big{]}\text{ for }\forall i\in\{1,2\}.$ (3)
Where $y=r+\gamma\min\limits_{j=1,2}{Q_{\bar{\theta_{j}}}}(s^{\prime},\pi_{\bar{\phi}}(s^{\prime}))$, $\bar{\phi}$ denotes the target network parameters. The actor is updated according to the Deterministic Policy Gradient (Fujimoto et al., 2018) At each training step $t$, the agent first randomly samples a batch of state transitions from the buffer and calculates the loss gradient as described above. It is worth noting that the proposed method in our paper is a plug-in and does not affect the training process of the reinforcement learning algorithm itself.
@@
DrQ is an algorithm based on maximum entropy reinforcement learning. Its training method are the same as SAC (Haarnoja et al., 2018). The soft value function is trained to minimize the squared residual error:
$J_{V}(\phi)=\mathbb{E}_{s\sim D}\left[\frac{1}{2}(V_{\phi}(s)-Q_{\theta}(s,a)+\log\pi_{\psi}(a|s))\right]$ (4)
where $a$ is sampled according to the current policy instead of the replay buffer. The soft Q-function parameters can be trained to minimize the soft Bellman residual: $J_{Q}(\theta)=\mathbb{E}_{(s,a)\sim D}[\frac{1}{2}(Q_{\theta}(s,a)-\hat{Q}(s,a))^{2}]$. Then the policy network can be learned by directly minimizing the expected KL divergence
$J_{\pi}(\psi)=\mathbb{E}_{s\sim D}[D_{KL}(\pi_{\phi}(\cdot|s)||\frac{\exp{Q(s)}}{Z(s)})]$ (5)
Appendix C Experimental Details
C.1 Network Structure
@@
The normal size of the TD3 network we used is the official architecture, and the detailed setting is shown in Tab 1. Besides, the network we used for Scale-up experiments incrementally adds the depth of both the actor and critic. For the 1.8$\times$ agent, we add one MLP as a hidden layer in two critic networks in front of the output head and one MLP layer as the embedding layer at the front of the actor. And the hidden dim is also increased to 256. And so on, we follow the above approach to grow the model by an equal amount each time.
@@
Fully Connected (128, action dim) and (128, 1)
@@ C.2 Implementation Details
Our codes are implemented with Python 3.8 and Torch 1.12.1. All experiments were run on NVIDIA GeForce GTX 3090 GPUs. Each single training trial ranges from 6 hours to 21 hours, depending on the algorithms and environments, e.g. DrQ spends more time than TD3 to handle the image input and DMC needs more time than OpenAI mujoco for rendering. Our TD3 is implemented with reference to github.com/sfujim/TD3 (TD3 source-code). The hyper-parameters for TD3 are presented in Table 2. Notably, for all OpenAI mujoco experiments, we use the raw state and reward from the environment and no normalization or scaling is used. An exploration noise sampled from $N(0,0.1)$ (Lillicrap, 2015) is added to all baseline methods when selecting an action. The discounted factor is ${0.95,0.99}$ and we use Adam Optimizer (Kingma, 2014) for all algorithms. Tab.2 shows the common hyperparameters of TD3 used in all our experiments. Following (Elsayed et al., 2024), We start applying weight clipping on both input and output weights of a neuron once it becomes part of the topology. To ensure a fair comparison and reproducibility of the expected performance of all baselines, we employ the recommended hyperparameter setting in Elsayed et al. (2024) for all baselines, with the clipping parameter $\kappa=3$ for MuJoCo tasks and $\kappa=2$ for image-input tasks. In our method, dormant neurons are randomly reinitialized after pruning, preparing them to rejoin the network topology. When they are re-selected, weight clipping is applied to the corresponding neurons. Weight clipping is a well-established, straightforward, and generic implementation technique (that operates independently of core algorithm design and specific algorithm choices, rather than a new algorithmic contribution here), which is a standard practice that we apply consistently across all implemented methods. Weight clipping is activated after the network initialization and operates throughout the training process. Additionally, we introduce layer normalization after the hidden layers of both the critic and actor networks to establish the LN-TD3 baseline for comparison. Notably, other methods do not incorporate normalization mechanisms to ensure a fair comparison. We reproduce Plasticity Injection Following the guidelines from the official paper (Nikishin et al., 2024). We utilize the default settings of the official ReDo codebase for our comparative analysis https://github.com/google/dopamine/tree/master/dopamine/labs/redo. For Reset-TD3, after a lot of empirical debugging, we found that parameter reset every $90,000$ step is a good setting, and fixed it as a hyperparameter. We suggest using seed 5 to reproduce the learning curve in Figure 5.
@@
Representation Model Learning Rate None None None $1e^{-4}$ $5e^{-3}$ $5e^{-3}$
@@
Table 2: A comparison of common hyperparameter choices of algorithms. We use ‘None’ to denote the ‘not applicable’ situation.